OpenStudy (anonymous):

A recipe that makes 7 servings calls for 1/2 cup of apple juice and cup of cranberry juice. How many cups of juice are there per serving? A. 1/6 B.3/5 C.1 1/16 D.8 1/6

OpenStudy (texaschic101):

how much cranberry juice ?

OpenStudy (anonymous):

oh sorry 2/3

OpenStudy (texaschic101):

add the juices 1/2 + 2/3 = 3/6 + 4/6 = 7/6 so there is 7/6 cups of juice for 7 servings. To find out how many are for each serving, we will use a proportion. (7/6) / 7 = x/1 --- (7/6 cups of j to 7 serv = x cups of j to 1 serv cross multiply 7x = 7/6 -- divide by 7 x = (7/6) / 7 x = 7/6 * 1/7 x = 7/42 which reduces to 1/6 cups per serving check.. 1/6 * 7 = 7/6 7/6 = 7/6 (correct) answer is 1/6 cups of juice per serving

OpenStudy (anonymous):

thanks dude

OpenStudy (texaschic101):

lol...no problem :)

OpenStudy (anonymous):

I have I few more can you help me

OpenStudy (texaschic101):

what ya got ?

OpenStudy (anonymous):

A recipe that makes 6 servings calls for 1/3 cup of orange juice and 1/2 cup of lemonade. Arlene modifies the recipe so that it can serve 9. How many cups of orange juice does Arlene need?

OpenStudy (texaschic101):

6/(1/3) = 9/x -- (6 serv to 1/3 o.j. = 9 serv to x cups o.j.) cross multiply 6x = (1/3)(9) 6x = 3 x = 3/6 which reduces to 1/2 cups of o.j. for 9 servings answer is 1/2 cups of orange juice for 9 servings

OpenStudy (anonymous):

A recipe for potting soil that makes enough to fill 3 bags calls for 5 1/2 lb of compost. David wants to modify the recipe so that it fills 5 bags. How many more pounds of compost will David need for the modified recipe than the original recipe calls for?

OpenStudy (texaschic101):

5 1/2 = 11/2 3/(11/2) = 5/x -- (3 bags to 11/2 lbs = 5 bags to x lbs) cross multiply 3x = (11/2)(5) 3x = 55/2 x = (55/2) / 3 x = 55/2 * 1/3 x = 55/6 which equals 9 1/6 lbs so he needs 9 1/6 (or 55/6) lbs and he has 5 1/2 (or 11/2) lbs 55/6 - 11/2 = 55/6 - 33/6 = 22/6 which equals 3 2/3 lbs he needs 3 2/3 more lbs

OpenStudy (anonymous):

k one more